OpenStudy (anonymous):

What are the exact solutions of x^2 - 5x - 1 = 0?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Thats not gonna help me, as ive said on the chat, I am doing an exam and neeed ANSWERS, not tips, but thank you for the tip tho, im very frustrated at the moment so if i seem like a jerk theres your reason

OpenStudy (lalaly):

do u know how to use the quadratic formula

OpenStudy (anonymous):

no. not reall, my profession im goin in doesnt require thay

OpenStudy (lalaly):

this equation cannot be factored, so u use the quadratic equation for \[ax^2+bx+c\] which is \[\frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] here a=1 b=-5 c=-1